OpenStudy (anonymous):

4x – 7y = –9 x + 3y = 12

OpenStudy (anonymous):

4x - 7y = -9 -------- eqn (1) x + 3y = 12--------eqn (2) Make x the subject for eqn (2) x = 12 - 3y --------eqn (3) Put the answer of x in eqn (3) into anywhere you see x in eqn (1) 4(12 - 3y) -7y = -9 Expand the bracket and solve for y 48 - 12y - 7y = -9 48 - 19y = -9 -19y = -9 - 48 -19y = -57 Divide both sides by -19 to make y stand alone \[y = \frac{-57}{-19}\] y = 3 Put the answer for y (y = 3) into eqn (2) above and solve for x x + 3y = 12 x + 3(3) = 12 x + 9 = 12 Subtract 9 from both sides to allow x to stand alone x = 12 - 9 x = 3 Hence, x = 3 and y = 3